From each of the following pairs of substances, use data in Appendix E in the textbook to choose the one that is the stronger oxidizing agent.

Match the words in the left column to the appropriate blanks in the sentences on the right. Make certain each sentence is complete before submitting your answer.

ResetHelp

  • HClO(aq)
  • O3(g)
  • H2O2(aq)
  • Zn2+(aq)
  • Fe2+(aq)
  • Cl2(g)
  • F2(g)
  • I2(s)
  • Given Zn2+(aq) or Fe2+(aq) , the stronger oxidizing agent is

    .GivenH2O2(aq)orO3(g), the stronger oxidizing agent is blank.

  • .GivenCl2(g)orHClO(aq), the stronger oxidizing agent is blank.
  • .GivenF2(g)orI2(s), the stronger oxidizing agent is blank.
Answer #1

Higher the value of standard reduction potential of a compound, it will have strong tendency to get reduced and oxidize the other compound. The compound with higher reduction potential will be stronger oxidizing agent.

Zn2+(aq) + 2e-Zn(s). E° = -0.76V

Fe2+(aq) +2e-Fe(s) . E° = -0.44V

(1) the stronger oxidizing agent is Fe2+(aq). (Answer)

H2O2 +2H+ + 2e- O2 + 2H2O . E° = 1.78V

O3 + 2H+ + 2e- O2 + H2O. E° = 2.07V

(2)the stronger oxidizing agent is O3(g). (Answer)

Cl2(g) + 2e-2Cl-(aq). E° = 1.36V

2HClO(aq) + 2H+ + 2e- Cl2 + 2H2O. E° = 1.63V

(3) the stronger oxidizing agent is HClO(aq) . (Answer)

F2(g) + 2e-2F-(aq). E° = 2.87V

I2(s) + 2e- 2I-(aq). E° = 0.54V

(4) the stronger oxidising agent is F2(g) . (Answer)